velmalovesshaggy145:

Maximum possible zeroes for the polynomial f(x)=-2x^8+2x^5-x^4+5x^2-9

B1acK1TT3n:

hmm i dont think i know this one tbh

Swirxy:

In order to isolate the variable in this linear equation, we need to get rid of the coefficient that multiplies it. This can be accomplished if both sides are divided by x .

velmalovesshaggy145:

-2x^7+2x^4-x^3+5x-9?

velmalovesshaggy145:

So 3 zeroes, since the degree is a zero?

velmalovesshaggy145:

ax^n-1 for 2x^4, which is 2x^4-1=2x^3 = 3 zeroes

velmalovesshaggy145:

@swirxy Thanks for the help

Swirxy:

Yw

surjithayer:

as it is of degree 8. possible number of zeros=8 f(x)=-2x^8+2x^5-x^4+5x^2-9 signs of f(x)=- + - + -9 no. of sign changes =4 so there are at the most 4 positive zeros. f(-x)=-2x^8-2x^5-x^4+5x^2-9 signs of f(-x)=- - - + - no. of sign changes in f(-x)=2 Hence there are two negative zeros.